OpenStudy (iwanttogotostanford):

PLZ HELP W/ MATH: Multiply: (3X -5)(2X+ 5)

OpenStudy (alexandervonhumboldt2):

FOIL: (a+b)*(c+d)=ac+ad+bc+bd (3x-5)*(2x+5)=3x*2x+5*3x-5*2x-5*5=6x^2+15x-10x-25 now combine like terms.

pooja195 (pooja195):



OpenStudy (iwanttogotostanford):

ok so i get 6x^2+15x-10x-25?

pooja195 (pooja195):

\[~\rm~(3x)(2x)+(3x)(5)+(−5)(2x)+(−5)(5)\]

pooja195 (pooja195):

Try again.

OpenStudy (alexandervonhumboldt2):

@pooja195 he is absolutely correct , why you said `try again`?

pooja195 (pooja195):